# Rescuing 331 bileptons from the Landau pole

### v1

@@ 4 Extending the scalar sector
The light scalar triplet $\rho$ is neutral under $U_{X}(1)$ (see Table 2), hence the addition of an arbitrary number of $\rho$-like triplets leaves $b_{X}$ invariant. This argument does not hold for the other weak $b_{i}$-coefficients. We add to the minimal 331 model an arbitrary number of $\rho$-like scalar triplets $N_{\rho}$ which correspond to extra Higgs $SU_{L}(2)$ doublets and contribute to the running between $\mu_{EW}$ and $\mu_{331}$. The $b_{i}$-parameters responsible of the running couplings below $\mu_{331}$ become (see eqs. (18) and (20)):
@@
Summarizing, the addition of $\rho$-like scalar triplets makes it possible to push $\mu^{\rm{max}}_{331}$ to higher energies, which allows accommodating a wider range of mass values for the bileptons and enlarging the energy range of validity (above $\mu_{331}$) of 331 extensions. We highlight that the case with $N_{H}=2+N_{\rho}=6$ can arise from an $E_{6}$ inspired framework. Indeed the 331 model can be embedded into $E_{6}$ as (see [27])
$E_{6}\to SU(2)\times SU(6)\to SU_{c}(3)\times SU_{L}(3)\times U(1)\,.$ (30)
@@
There are other ways to extend the scalar sector. So far, we have considered one, two and many light Higgs $SU_{L}(2)$ doublets that are embedded into $SU_{L}(3)$ triplets. The triplets represent minimal configurations for the scalar sector. Maintaining the natural assumption that fermions transform either as the representation ${\bf 3}$ or ${\bf\bar{3}}$ under $SU_{L}(3)$, gauge invariance also allows the possibility of scalar sextets (see e.g. [16, 28]). It has to be noted that, should we choose to evade this assumption, e.g. including octet leptons [29] or scalar leptoquarks assigned to the representation ${\bf 8}$ of $SU_{L}(3)$ [12], a wider choice of representations for the scalar sector would be available, and thus further ways to move the Landau pole.
Let us now extend the scalar sector of the minimal 331 with a sextet $S$ of zero $X$-charge, which contributes to break spontaneously the 331 symmetry. Its presence allows the addition of a renormalizable [^3] In some 331 model, neutrino gets a Majorana mass by means of radiative contribution induced by triplet scalar representation beside extra gauge bosons [1]. Majorana neutrino mass term [23]
@@
where $L_{i}$ ($i=1,2,3$) are the left-handed lepton $SU_{L}(3)$-triplets. Since the sextet representation of $SU(3)$ decomposes into representations of the $SU(2)$ subgroup as $\mathbf{6}\rightarrow\mathbf{3}\oplus\mathbf{2}\oplus\mathbf{1}$, below $\mu_{331}$ we have one triplet and one iso-doublet and one scalar. In order to give a Majorana mass to neutrinos in the SM through the type II seesaw mechanism, one needs a $SU_{L}(2)$ scalar triplet to couple with lepton $SU_{L}(2)$ doublets. Following [12] we assume $SU_{L}(2)$ doublet and triplet contained in the sextet $S$ to be light, while the singlet heavy. The $b_{i}$-coefficient of eq.(18) are modified as
@@
while $b_{X}$ is unchanged. Using these $b_{i}$-parameters we obtain a value for the scale $\mu_{\rm max}\sim 5500$ GeV as shown in figure (4) with the dashed line. However if we assume that the new fields belonging to the sextet participate to the running staring from a scale $\mu\sim 600$ GeV, then $\mu_{\rm max}\sim 4630$ GeV, see figure (4) continuous line.
Figure 4: Running of $\alpha_{X}^{-1}$ in the sextet extension for different values of the mass of the doublet and triplet: electroweak scale ($dashed$ $line$), $600$ GeV (continuous line). .
@@
As far as we know, experimental constraints on the sequential fourth family extending minimal 331 models are not available yet–some model-dependent bounds are given in [36, 37]. Some insight can be gained by considering limits from vector-like quark searches. In particular we have $m_{\tau^{\prime}}>700$ GeV [38] and $m_{b^{\prime}}>1390\div 1570$ GeV[39] from the CMS Collaboration (the range is due to the variation of branching ratios) and $m_{t^{\prime}}>715\div 950$ GeV [40] and $m_{b^{\prime}}>1000\div 2000$ GeV [41] from the ATLAS Collaboration. A combined CMS and ATLAS study [42] reports exclusion limits for masses below about 1500 GeV of both $b^{\prime}$ and $t^{\prime}$. We remark again that the previous limits cannot be directly applied to our case. In CMS and ATLAS analyses the lepton universality is not required, while it is mandatory in a sequential fourth family model.
For the sake of simplicity, we assume $b^{\prime},t^{\prime},T^{2}$ to have the same mass $m_{\rm NP}$ and we consider e.g. the cases where $m_{\rm NP}=400,600,800,1000,1500$ GeV. We proceed in two steps: we run $\alpha^{-1}_{X}(\mu)$ from the electroweak breaking scale to $m_{\rm NP}$ using the $b$-parameters (20). Then we run it above $m_{\rm NP}$ using the $b$-parameters with:
$b_{Y}$ $=$ $\frac{83}{9},\qquad b_{2L}=-\frac{5}{3}\qquad\text{for $\mu<\mu_{331}$}\,,$ (34)
$b_{X}$ $=$ $\frac{134}{3},\qquad b_{3L}=-\frac{31}{6}\qquad\text{for $\mu>\mu_{331}$}\,.$ (35)
By using these parameters we determine the dependence of $\alpha_{X}^{-1}$ on $\mu_{331}$ for different values of the mass scale $m_{\rm NP}$. This is shown in left plot of figure (5).
Figure 5: Running of $\alpha_{X}^{-1}$ with the energy scale $\mu_{331}$. (Left plot) 4th family model for different values of the mass scale $m_{\rm NP}$: 1500 GeV (blue continuous), 1000 GeV (red continuous), 800 GeV (dashed), 600 GeV (dot-dashed) and 400 GeV (dotted). (Right plot) 4th family model + sextet with $m_{\rm NP}\sim 1500$ GeV.
Given the experimental bounds previously mentioned, a significative case is $m_{\text{NP}}\sim 1500$ GeV, where we find $\mu^{\text{max}}_{331}\sim 4400$ GeV.
As a final step we also include the scalar sextet (see section 4), obtaining $\mu^{\text{max}}_{331}=4919$ GeV as shown in right plot of figure (5).
We compare the minimal 331 model with the different extensions examined in this work and report the values of $b$-parameters and $\mu^{\rm{max}}_{331}$ in Table 1. We also report the bileptons’ masses estimated at $\mu^{\rm{max}}_{331}$ according to eq. (13).
@@
$\rho$-like triplets ($N_{H}=6$) 47/6 13/6 5500 1788
4th fam. ($\text{m}_{\text{NP}}\sim 1500$ GeV) 83/9 5/3 4400 1430
4th fam. + sextet ($\text{m}_{\text{NP}}\sim 1500$ GeV) 187/18 5/6 4919 1599
Table 1: Summary of $b$-parameters and corresponding values of the Landau pole for different 331 models and expected bilepton mass.
Moreover we add as a last example the case where we include the sextet to the four family case.
6 Conclusions
331 extensions of the Standard Model are characterized by a free parameter $\beta$. By setting $\beta=\sqrt{3}$, 331 models predict doubly charged gauge bosons that mainly decay into pairs of same-sign leptons and are therefore dubbed bileptons. Their production and decay yield distinctive signatures that enable direct searches at colliders.
@@ 6 Conclusions
We begin by analyzing the running of the gauge coupling constants in the minimal 331 model, considering its matter content and the associated scales of spontaneous symmetry breaking for both the 331 and electroweak sectors. We find that, to comply with experimental constraints, this model must behave as a Two Higgs Doublet Model in the energy range between $\mu_{\rm{EW}}$ and $\mu_{331}$, with both Higgs bosons acquiring mass within this range. Unfortunately, we also find that it is not possible to shift the Landau pole by varying the free parameters, namely $\mu_{331}$ and the masses of the extra quarks. This remains true even in the so-called next-to-minimal 331 model, which includes extra lepton fermions.
Viable solutions require extending the scalar sector and/or adding a sequential fourth family of non-standard quarks and leptons. The running below $\mu_{331}$ of the $U_{Y}(1)$ and $SU(2)_{L}$ coupling constants changes upon supplementing the minimal model with $\rho$-like Higgs triplets (corresponding to extra Higgs $SU_{L}(2)$ doublets). Although this procedure is constrained by bounds on the behavior of the $SU_{L}(2)$ coupling constant, the addition of scalar triplets of this kind extends the valid energy range, which in turn allows for a wider spectrum of bilepton masses, $m_{Y}\sim 1788$ GeV if $N_{H}=6$. Notably, the case with $N_{H}=6$ can originate from an $E_{6}$-inspired framework.
Another scalar sector extension analyzed in this work include an $SU_{L}(3)$ sextet. This case, which allows the introduction of a mass term for Majorana neutrinos, predicts $m_{Y}\sim 1472$ GeV.
Finally, we have studied an extension of the fermionic sector involving a sequential fourth family with two Higgs doublets. By making conservative assumptions on the mass of the fourth family based on recent experimental results, one manages to shift the Landau pole, and get $m_{Y}\sim 1472$ GeV with an intermediate mass scale of 1500 GeV. For the sake of completeness, we also study this model including the scalar sextet obtaining the higher value $m_{Y}\sim 1599$ GeV at the same intermediate scale. We summarize the results of the benchmark models in Table 1. We observe that extending the minimal 331 models increases the Landau pole, yet this increase remains within the range of experimental values attainable at LHC in the near future.
@@ Acknowledgements
Scalar $SU_{L}(3)$ $U_{X}(1)$
$\chi$ 3 $+1$
$\rho$ 3 0
$\eta$ 3 $-1$
Table 3: Matter content of the 331 model with 4th family: we use the notation of Ref. [31].

### v2

@@ IV Extending the scalar sector
The light scalar triplet $\rho$ is neutral under $U_{X}(1)$ (see Table 3), hence the addition of an arbitrary number of $\rho$-like triplets leaves $b_{X}$ invariant. This argument does not hold for the other weak $b_{i}$-coefficients. In the minimal 331 model, we add an arbitrary number of $\rho$-like scalar triplets $N_{\rho}$ which correspond to extra Higgs $SU_{L}(2)$ doublets and contribute to the running between $\mu_{EW}$ and $\mu_{331}$. The $b_{i}$-parameters responsible of the running couplings below $\mu_{331}$ become (see eqs. (18) and (20)):
@@
Summarizing, the addition of $\rho$-like scalar triplets makes it possible to push $\mu^{\rm{max}}_{331}$ to higher energies, which allows accommodating a wider range of mass values for the bileptons and enlarging the energy range of validity (above $\mu_{331}$) of 331 extensions. We highlight that the case with $N_{H}=2+N_{\rho}=6$ can arise from an $E_{6}$ inspired framework. Indeed the 331 model can be embedded into $E_{6}$ as (see e.g.  [28])
$E_{6}\to SU(2)\times$ $SU(6)\to$
$\to$ $SU_{c}(3)\times SU_{L}(3)\times U(1)\,.$ (30)
@@
There are other ways to extend the scalar sector. So far, we have considered one, two and many light Higgs $SU_{L}(2)$ doublets that are embedded into $SU_{L}(3)$ triplets. The triplets represent minimal configurations for the scalar sector. Maintaining the natural assumption that fermions transform either as the representation ${\bf 3}$ or ${\bf\bar{3}}$ under $SU_{L}(3)$, gauge invariance also allows the possibility of scalar sextets (see e.g. [15, 29, 30]). It has to be noted that, should we choose to evade this assumption, e.g. including octet leptons [31] or scalar leptoquarks assigned to the representation ${\bf 8}$ of $SU_{L}(3)$ [12], a wider choice of representations for the scalar sector would be available, and thus further ways to move the Landau pole.
Let us now extend the scalar sector of the minimal 331 with a sextet $S$ of zero $X$-charge, which contributes to break spontaneously the 331 symmetry. Its presence allows the addition of a renormalizable  [^3] Let us remark that there are 331 models where neutrino gets a Majorana mass by means of radiative contribution induced by triplet scalar representation, beside extra gauge bosons [1]. Majorana neutrino mass term [24]
@@
where $L_{i}$ ($i=1,2,3$) are the left-handed lepton $SU_{L}(3)$-triplets. Since the sextet representation of $SU(3)$ decomposes into representations of the $SU(2)$ subgroup as $\mathbf{6}\rightarrow\mathbf{3}\oplus\mathbf{2}\oplus\mathbf{1}$, below $\mu_{331}$ we have one triplet and one iso-doublet and one scalar. In order to give a Majorana mass to neutrinos in the SM through the type II seesaw mechanism, one needs a $SU_{L}(2)$ scalar triplet to couple with lepton $SU_{L}(2)$ doublets. Following [12] we assume $SU_{L}(2)$ doublet and triplet contained in the sextet $S$ to be light, while the singlet heavy. The $b_{i}$-coefficients of eq.(18) are modified as
@@
while $b_{X}$ is unchanged. Using these $b_{i}$-parameters we obtain a value for the scale $\mu^{\rm max}_{331}\sim 5500$ GeV as shown in figure (4) with the dashed line. However if we assume that the new fields belonging to the sextet participate to the running staring from an intermediate scale, let us say $\mu\sim 600$ GeV, then $\mu^{\rm max}_{331}\sim 4630$ GeV, see figure (4) continuous line.
Figure 4: Running of $\alpha_{X}^{-1}$ in the sextet extension for different values of the mass of the doublet and triplet: electroweak scale ($dashed$ $line$), 600600600 GeV (continuous line).
@@
As far as we know, experimental constraints on the sequential fourth family extending minimal 331 models are not available yet. Model-dependent bounds in different frameworks are given in [39, 40]. Some insight can be gained by considering limits from vector-like quark searches. In particular we have $m_{\tau^{\prime}}>700$ GeV [41] and $m_{b^{\prime}}>1390\div 1570$ GeV[42] from the CMS Collaboration (the range is due to the variation of branching ratios) and $m_{t^{\prime}}>715\div 950$ GeV [43] and $m_{b^{\prime}}>1000\div 2000$ GeV [44] from the ATLAS Collaboration. A combined CMS and ATLAS study [45] reports exclusion limits for masses below about 1500 GeV of both $b^{\prime}$ and $t^{\prime}$. We remark again that the previous limits cannot be directly applied to our case. In CMS and ATLAS analyses the lepton universality is not required, while it is mandatory in a sequential fourth family model.
We note that requiring unitary at very high energies in $4th$ generation models, suggests a bound on the BSM fermion masses that is $m_{f}^{\prime}\lesssim 500$ GeV [46]. However as reported in [35] this bound should not be viewed as a strict limit but simply as an indication that strong dynamics takes place. Moreover simulation of a strong Yukawa coupling regime on the lattice found $m_{f}^{\prime}\lesssim 700$ GeV [47]. Therefore hereafter we will consider the special scale $m_{\rm NP}\sim 700$ GeV as a benchmark case.
For the sake of simplicity, we assume $b^{\prime},t^{\prime},T^{2}$ to have the same mass $m_{\rm NP}$ and we consider e.g. the cases where $m_{\rm NP}$=500, 700, 1000, 1500 GeV. We proceed in two steps: we run $\alpha^{-1}_{X}(\mu)$ from the electroweak breaking scale to $m_{\rm NP}$ using the $b_{i}$-parameters (20). Then we run it above $m_{\rm NP}$ using the following $b_{i}$-parameters:
$b_{Y}$ $=$ $\frac{83}{9},\quad b_{2L}=-\frac{5}{3}\qquad\text{for $\mu<\mu_{331}$},$ (34)
$b_{X}$ $=$ $\frac{134}{3},\quad b_{3L}=-\frac{31}{6}\quad\text{for $\mu>\mu_{331}$}.$ (35)
By using these parameters we determine the dependence of $\alpha_{X}^{-1}$ on $\mu_{331}$ for different values of the mass scale $m_{\rm NP}$. This is shown in the upper plot of figure (5).
Figure 5: Running of $\alpha_{X}^{-1}$ with the energy scale $\mu_{331}$. (Top plot) 4th family model for different values of the mass scale $m_{\rm NP}$: 1500 GeV (dotted), 1000 GeV (dot-dashed), 700 GeV (dashed), 500 GeV (continue). (Bottom plot) 4th family model + sextet with $m_{\rm NP}\sim 700$ GeV.
In the benchmark case $m_{\text{NP}}\sim 700$ GeV, we find $\mu^{\text{max}}_{331}\sim 4840$ GeV.
As a final step we also include the scalar sextet (see section IV), obtaining $\mu^{\text{max}}_{331}=5912$ GeV as shown in the lower plot of figure (5).
In Table 2, we compare the different extensions of the minimal 331 models we examined. We report the values of $b_{i}$-parameters in the energy range between the electroweak and the 331 symmetry breaking scales, the values of $\mu^{\rm{max}}_{331}$ and the bileptons’ masses estimated at $\mu^{\rm{max}}_{331}$, according to eq. (13).
@@
$N_{H}=6$ 47/6 13/6 5500 1788
[…]
Figure 6: Benchmark models running report
VI Conclusions
The 331 extensions of the Standard Model are characterized by a free parameter $\beta$. By setting $\beta=\sqrt{3}$, 331 models predict doubly charged gauge bosons that mainly decay into pairs of same-sign leptons and are therefore dubbed bileptons. Their production and decay yield distinctive signatures that enable direct searches at colliders.
@@ VI Conclusions
We begin by analyzing the running of the gauge coupling constants in the minimal 331 model, considering its matter content and the associated scales of spontaneous symmetry breaking for both the 331 and electroweak sectors. We find that, to comply with experimental constraints, this model must behave as a Two Higgs Doublet Model in the energy range between $\mu_{\rm{EW}}$ and $\mu_{331}$, with both Higgs bosons acquiring mass within this range. Unfortunately, we also find that it is not possible to shift the Landau pole by varying the free parameters, namely $\mu_{331}$ and the masses of the exotic quarks. This remains true even in the so-called next-to-minimal 331 model, which includes additional non-standard leptons.
Viable solutions require extending the scalar sector and/or adding a sequential fourth family of non-standard quarks and leptons. The running below $\mu_{331}$ of the $U_{Y}(1)$ and $SU(2)_{L}$ coupling constants changes upon supplementing the minimal model with $\rho$-like Higgs triplets (corresponding to extra Higgs $SU_{L}(2)$ doublets). Although this procedure is constrained by bounds on the behavior of the $SU_{L}(2)$ coupling constant, the addition of scalar triplets of this kind extends the valid energy range, which in turn allows for a wider spectrum of bilepton masses, e.g. $m_{Y}\sim 1788$ GeV if $N_{H}=6$. Notably, the case with $N_{H}=6$ can originate from an $E_{6}$-inspired framework.
Another scalar sector extension analyzed in this work includes an $SU_{L}(3)$ sextet. This case, which allows the introduction of a mass term for Majorana neutrinos, predicts $m_{Y}\sim 1472$ GeV.
Finally, we have studied an extension of the fermionic sector involving a sequential fourth family with two Higgs doublets. By making conservative assumptions on the mass of the fourth family based on recent experimental results, one manages to shift the Landau pole, and get $m_{Y}\sim 1573$ GeV with an intermediate mass scale of 700 GeV. For the sake of completeness, we also study this model including the scalar sextet, obtaining the higher value $m_{Y}\sim 1921$ GeV at the same intermediate scale.
We summarize our results in Table 2 and Figure 6. Extending the minimal 331 models increases the Landau pole, yet this increase remains within the range of experimental values attainable at LHC in the near future.
@@ Acknowledgements
Table 4: Matter content of the 331 model with 4th family.